OpenStudy (anonymous):

A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6)?

OpenStudy (freckles):

let n be 1 for now f(1+1)=f(1)-8 hmmm 1+1=2 so we have f(2)=f(1)-8 hmm f(1) is given as 100 f(2)=100-8=92 f(2)=92 see if you can find f(3) now

OpenStudy (freckles):

you will have to do this a few more times because to find f(6) we need to know f(5) and the find f(5) we need to know f(4) and to find f(4) we need to know f(3)

OpenStudy (freckles):

so noticed i replace n with 1 in the beginnning replace n's with 2 instead then keep one upping it

OpenStudy (freckles):

like this replace n with 2 f(2+1)=f(2)-8 this will give you f(3)=f(2)-8 which I think you can find now